A 150-kg roller coaster car moving with a velocity of 4 m/s to the right hits a 125-kg car that is standing still. If the two cars hit and stick together, what is their final velocity?
1
Expert's answer
2019-06-11T10:10:26-0400

To determine the velocity of the two cars that stuck together in a collision, we use the law of conservation of momentum: the total momentum of the system before and after the collision remains the same

"{{\vec{p}}_{1}}+{{\vec{p}}_{2}}=\text{constant}"


Recall that the momentum of an object is the product of its mass m and velocity "\vec{v}"


"\vec{p}=m\vec{v}"


Then the law of conservation of momentum is written as


"{{m}_{1}}{{\vec{v}}_{1}}+{{m}_{2}}{{\vec{v}}_{2}}={{m}_{1}}{{\vec{u}}_{1}}+{{m}_{2}}{{\vec{u}}_{2}}"


where "{{m}_{1}}=150\,kg,\,\,{{m}_{2}}=125\,kg" are the masses of the first and second cars, "{{\vec{v}}_{1}}=4\,m/s,\,{{\vec{v}}_{2}}=0\,m/s" are their velocities before the collision and "{{\vec{u}}_{1}}={{\vec{u}}_{2}}=\vec{u}" are their velocities after the collision (final velocity). Also define the direction of the initial velocity vector to be the x-direction. Now this equation becomes


"{{m}_{1}}{{v}_{1}}\cdot \vec{i}+0=\left( {{m}_{1}}+{{m}_{2}} \right)\vec{u}"


Find the final speed


"\vec{u}=\frac{{{m}_{1}}{{v}_{1}}}{{{m}_{1}}+{{m}_{2}}}\vec{i}"


Substituting the given numbers:


"\vec{u}=\left( \frac{150\,kg\cdot 4\,m/s}{150\,kg+125\,kg} \right)\vec{i}\approx \left( 2.2\,m/s \right)\vec{i}"